OpenStudy (anonymous):

4 people can paint 6 wall is 30 minutes. How many minutes will it take 7 people to paint 9 walls???

OpenStudy (anonymous):

Try to understand the concept from here. http://openstudy.com/users/siddhantsharan#/updates/4f5a2f42e4b0636d8905cb79

OpenStudy (anonymous):

4 ppl ----6 walls ----30 mins. 1 per-----6/4 wall ---30mins 1 per------1wall ---30*4/6 mins 7ppl ----9 walls 1 pr needs----- 9/7walls Thereore time = 9/7*30*4/6 25.7 mins

OpenStudy (phi):

I think of it as people*time = work and you need a certain amount of work to finish a job. so as a ration \[ \frac{4*30}{6}=\frac{7t}{9} \] now solve for t by multiplying both sides by 9/7

OpenStudy (phi):

*ratio